if xyz=4 and y^2z=5, what is the value of x/y?
Answer by ovie27(16)


xyz=4 and y^2z=5
So, we can solve for z from the second equation:
z = 5/y^2
and we can plug this into the first one:
xy(5/y^2)=4
Cross out the y:
x5/y = 4
5x = 4y
x = 4y/5
divide by y:
x/y = 4/5